Taussig Technologies Corporation (TTC) has been growing at a rate of 20% per year in recent years. This same growth rate is expected to last for another 2 years, then to decline to gn = 6%.

a.   If D0 = $1.60 and rs = 10%, what is TTC's stock worth today? What are its expected dividend
      and capital gains yields at this time, that is, during Year 1?
Answer #1

a. D0 = $1.60

D1 = D0 (1 + g s) = $1.6(1.20) = $1.92.

D2 = D0 (1 + g s)^2 = $1.60(1.20)^2 = $2.304.

D3 = D2 (1 + g n) = 2.304 (1.06) = 2.44224

P2 = D3 /(Ks – g n) =

P2 = 2.44224 / (0.10 – 0.06) = $61.06

P0 = PV of D1 + PV of D2 + PV of P2

= D1/ (1 + 0.10) + D2/ (1 + 0.10)^2 + P2 / ( 1 + 0.10)^2

= $1.92*0.909091 + $2.304 * 0.82645 + 61.06* 0.82645

= $ 54.1091

2.

Dividend yield = D1 / P0

Dividend yield = $ 1.92 / $ 54.1091

Dividend yield = 3.55%

3.

The capital gains yield can be calculated by simply subtracting the dividend yield from the total expected return

Cap. Gain yield = Expected return - Dividend yield

Cap. Gain yield = 10.0% - 3.55%

Cap. Gain yield = 6.45%